5(x-3)^2+4=7 how would you solve this I have no idea
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

5(x-3)^2+4=7
-------------
5(x-3)^2 = 3
----
(x-3)^2 = 3/5
----
x-3 = sqrt(3/5) or x-3 = -sqrt(3/5)
===
x = 3+sqrt(15)/5 or x = 30-sqrt(15)/5
==============================
Cheers,
Stan H.
===========

Answer by lynnlo(4176)

5(x-3)^2+4=7
5(x-3)^2+4-7=0
5(x-3)^2+3-=0
5(x-3)^2-3=0
5(x-3)^2-3=0
RESULT=============0=0
SOLUTION===========X=3,,,,X=2